OpenStudy (alex6799):

Use a graphing calculator to solve the equation in the interval from Round your answers to the nearest hundredth. A. 1.21, 4.35 B. 3.64 C. 1.21, 2.26, 3.31, 4.35, 5.40 D. 0.404, 1,452.5, 3.55, 4.59, 5.64

OpenStudy (alex6799):

\[3\tan \frac{ 1 }{ 3 }0=8\] \[0 \to 2\pi\]

OpenStudy (alex6799):

how do i get an answer from that?!

OpenStudy (alex6799):

ok?

OpenStudy (aihberkhan):

I know... this is a very confusing one... umm this is confusing me. I don't think it is best if I help you with this one, because i don't want to give you the wrong answer. So lets just ask someone else to help! :) I can help get somewhere here by tagging them ;)

OpenStudy (aihberkhan):

@DanJS Please help alex

OpenStudy (aihberkhan):

@imqwerty Can you help alex? :)

imqwerty (imqwerty):

okay you have this- \(\large 3 tan\Large \frac{1}{3} \theta=8\) and you need a solution in this range-> \([0,~2 \pi]\) okay 1st we will simplify it a bit :) \(\large 3 tan\Large \frac{1}{3} \theta=8\) \(\large tan\Large \frac{\theta}{3}=\frac{8}{3}\) \(\large tan\Large \frac{\theta}{3}=2.66\) lets say that \(\large \frac{\theta}{3}= \alpha\) since \(\theta\) ranges from \([0, ~2 \pi]\) \(\alpha\) ranges from \([0,~\large \frac{2 \pi}{3}]\) now we have-> \(\large \tan(\alpha)=2.66\) now we open up our graphing calculator and draw the graph for \(\tan(\alpha)\) and the point between [0, \(\frac{2 \pi}{3}]\) where the value is 2.66 is our \(\alpha\) now we have \(\alpha\) we just have to multiply it by 3 to get \(\theta\) because \(\large \frac{\theta}{3}= \alpha\)

OpenStudy (alex6799):

so C?

OpenStudy (alex6799):

@imqwerty

imqwerty (imqwerty):

no its not c

OpenStudy (alex6799):

ok, then it has to be B.......right?

imqwerty (imqwerty):

yea b is correct :)

OpenStudy (alex6799):

YAY

OpenStudy (alex6799):

i have two questions i need help with then im done....will/can either one of you help me with them really quick?